OpenStudy (anonymous):

"The 15-kg block A slides on the surface for which μk = 0.3. The block has a velocity v = 10 m/s when it is s = 4 m from the 10-kg block B. The unstretched spring has a stiffness k = 1000 N/m . Take e = 0.6. The coefficient of friction is the same for both blocks. Determine the maximum compression of the spring due to the collision." I got x(max) = 0.839 m, but it's saying that it's not quite right. Can anyone help me?

OpenStudy (anonymous):

Here's an image of the problem. If you would like, I can show how I arrived at my solution.

OpenStudy (johnweldon1993):

Absolutely! Lets see where the work went off

OpenStudy (anonymous):

Okay, but this is going to take a minute for me to type up I have it written down on paper here, so I'll need some patience. Please hold.

OpenStudy (johnweldon1993):

No problem!

OpenStudy (anonymous):

Okay, beginning with the free body diagrams;  Summing forces in the y direction for A gives N_A = m_A*g = (15 kg)(9.81 m/s) = 147.15N f_k = (coefficient kin. frict.)(N_A) = (0.3)(147.15) = 44.145 N. Using Conservation of Energy for A: T_1 + V_1 + (SIgma)U_1->2 = V_2 + T_2 ==> (1/2)m_A*v - s*(co. kin. frict.)(N_A) = (1/2)*m_A*(v_A)^2 ==> (1/2)(15 kg)(10 m/s)^2 - (4 m)(44.145 N) = (1/2)(15 kg)(v_A)^2 ==> v_A = 8.74 m/s You with me so far?